|
Inspection of side FEEs (2,6,4,8) and possible grounding improvements:
- all screws and nuts for copper connections were tightenedFee
- FEE 6 was isolated from the copper connection grounding FEE2 with insulating tape (attachment 1)
- A piece of tape was found below FEE6 (attachments 2,3) --> shall this be removed?
Bias and power on of FEES, preparation of two settings with lower thresholds:
- note: cannot calibrate ADC of aida15 even after trying multiple times (?)
- 500-keV thresholds settings saved as EXPERIMENTS/AIDA/2025Oct29-09.49.01
- 200-keV thresholds settings saved as EXPERIMENTS/AIDA/2025Oct29-09.55.31
|